# UniPLV: Towards Label-Efficient Open-World 3D Scene Understanding by Regional Visual Language Supervision

### v1

@@ Abstract
We present UniPLV, a powerful framework that unifies point clouds, images and text in a single learning paradigm for open-world 3D scene understanding. UniPLV employs the image modal as a bridge to co-embed 3D points with pre-aligned images and text in a shared feature space without requiring carefully crafted point cloud text pairs. To accomplish multi-modal alignment, we propose two key strategies:(i) logit and feature distillation modules between images and point clouds, and (ii) a vison-point matching module is given to explicitly correct the misalignment caused by points to pixels projection. To further improve the performance of our unified framework, we adopt four task-specific losses and a two-stage training strategy. Extensive experiments show that our method outperforms the state-of-the-art methods by an average of 15.6% and 14.8% for semantic segmentation over Base-Annotated and Annotation-Free tasks, respectively. The code will be released later.
1 Introduction
Figure 1: The matching example of 2D images and 3D point clouds. (a): Expected matching between images and point clouds. (b): Image semantic labels are assigned to the corresponding point clouds by projection, which has many false positive mismatched points. (c): The semantic predictions of new classes without any manual annotations by our method. (d): Semantic prediction of our method for all points of an input point cloud.
Open-world 3D scene understanding aims to obtain a model that can distinguish and recognize open-set objects and categories from 3D data, e.g. point clouds, without manual annotations. Traditional closed-set 3D understanding relying on manual annotations cannot accomplish the task of open-set recognition. Therefore, open-world 3D scene understanding is critical for real-world applications, such as autonomous driving, mobile robots, virtual reality, etc.
Recent success of visual language models(VLM) has provided opportunities for open-world 3D scene understanding. Frozen-VLM labeling-based methods [38, 22] transfer the open vocalbulary of VLM to 3D scenes for implementing open-world 3D scene understanding. These methods neglect region-level feature distillation learning and require resource-intensive feature extraction, fusion, and storage processes[40], thus hindering scalability to more advanced 3D networks and 3D scenes, which in turn makes it difficult to improve performance. Pairs-construction based methods [8, 7, 33] attempts to align point clouds and text like VLM training by constructing point clouds text pairs. RegionPLC [33] achieves the state-of-the-art (SOTA) performance, but requires a tedious process to produce a large number of point clouds text pairs for contrastive training. However, unlike VLM that can obtain large-scale text-image pairs from the Internet, point cloud text pairs cannot be easily obtained, resulting in performance limitations for pairs-construction based methods.
Fortunately, although the number of point clouds data is limited, point clouds typically appear in pairs with images, in which the calibration and time compensation are generally completed. A straightforward approach to achieve 3D open-world understanding involves initially applying 2D open-world models for labeling on images, followed by 3D-2D projecting to derive the corresponding semantic categories (refer 4.4 for details). However, as shown in Figure 1 (b), the projection between images and point clouds is inaccurate (due to factors such as vehicle motion bumps, calibration and compensation errors, etc.), causing numerous false positives and omissions in the projected 3D semantics. Given the spatially paired point clouds and images, along with feature-aligned images and text in visual language models, images can serve as a bridge in a unified learning framework, easing the challenge of aligning text with point clouds. To implement the unified open-world 3D understanding framework with limited 3D data, the following challenges must be addressed: (i) co-embedding the point cloud features with pre-aligned images and text in a shared space, despite limited data availability, (ii) correcting projection misalignments between point clouds and images, and (iii) stably training the unified multi-modal framework to achieve the excellent performance.
Specifically, we begin by constructing the image regional semantic labels by 2D foundation models to supervise the new categories during training. UniPLV integrates a point clouds understanding network and a pre-aligned visual language model, where the image branch is the bridge we need to use. With the image-text alignment and the spatial projection between point clouds and image, we design the classification logits distillation and the feature distillation between image and point clouds modalities to co-embed point clouds with text and images in the same feature space. We further propose a vison-point matching (VPM) learning module to refine point-pixel matching and correct point clouds text alignment. To address the challenges of novel classes being buired as long-tail categories and the gradient bias in multi-modal training, we introduce a two-stage training procedure and four task-specific weighted losses, including image-text alignment, point-text alignment, pixel-point matching, and distillation losses. The training begins with an initial stage focused on independently training the image branch for half of the total iterations, followed by a joint training stage for both image and point clouds branches with distinct loss weights to effectively balance their contributions. Simply providing text descriptions and point clouds, our method could obtain 3D semantic understanding of point clouds during inference.
In summary, our contributions can be summarized into four-fold:
@@ 1 Introduction
  • We propose a unified multi-modal contrastive learning framework for open world 3D scene understanding without constructing point clouds text pairs.
@@
  • We propose logit distillation, feature distillation, and vision-point matching modules to co-embed three modalities in a shared feature space.
@@
  • We propose a two-stage optimization strategy and four task-specific weighted losses to ensure the stable and efficient multi-modal training.
  • Our proposed method achieves the state-of-the-art performance for two open world tasks conducted on nuScenes, and is further evaluated on two additional competitive outdoor datasets, Waymo and SemanticKITTI.
2 Related works
3D Semantic Segmentation. Semantic segmentation for LIDAR point clouds can be categorized into three categories based on the form of modeling points, i.e., view-based, point-based, and voxel-based methods. View-based methods convert 3D point clouds to the range view [20, 16, 17] or the bird-eye[41] view to extract 2D features with loss of 3D geometric characteristics. Point-based methods [23, 14, 43, 30, 31] directly use points as inputs to the model and design algorithms to assemble contexts. Voxel-based approaches [6, 9, 42, 45, 10] subdivide the point cloud space into numerous voxel grids and process these voxel features using sparse convolution techniques to improve effectiveness. This paper uses MinkUNet [6], SparseUnet32 [9], and PTV3 [31] as the backbone network to verify the scalability and lightweight of the proposed framework, respectively.
@@ 2 Related works
Open-Vocabulary 2D Scene Understanding. The recent advances of large visual language models have expanded the ability to understand 2D open-world scenes. There are two main directions: CLIP-based and Grounding. CLIP-based methods usually replace the linear projection feature with the CLIP feature and use CLIP for feature alignment, such as GLEE[29], DetCLIP series[35, 36, 37], RegionCLIP[44], OWL-ViT[21]. The input of Grounding tasks is a picture and a corresponding description. Grounding tasks output object boxes at the corresponding positions in the image based on different descriptions [15, 39, 18, 26]. Given the success of 2D open world understanding, we select GLEE[29] and Grounding DINO[18, 26] as our 2D open set region label generators, respectively.
@@
Open-Vocabulary 3D Scene Understanding. Open-world 3D scene understanding aims to recognize objects that are not annotated. Early methods [19, 5, 32, 3] mainly achieve open-world understanding through feature differentiation or generation approaches. With the success of visual-language (VL) models, such as CLIP [24], numerous efforts [4, 28, 40, 22, 12] have emerged to transfer the VL knowledge to 3D scene understanding. Clip2Scene [4] utilizes frozen CLIP to obtain semantic labels of images, which are subsequently projected to guide a point cloud semantic segmentation. OpenMask3D [28] employs a 3D instance segmentation network to create 3D masks, which are projected to obtain 2D masks. These 2D masks are then fed into CLIP to extract visual features and match them with textual features, ultimately obtaining 3D semantics. Since CLIP is trained based on full image and text alignments, its capability to comprehend specific regions is limited. OpenScene [22] facilitates point cloud and text alignment by projecting the prediction results from a frozen 2D vision model and employing distillation between image and point clouds features. However, OpenScene requires resource-intensive feature extraction and fusion, and the image backbone is fixed during training, making it unable to be easily expanded to more advanced 3D networks and 3D scenes. RegionPLC [34] and PLA [7] achieve open world 3D scene understanding by constructing a large number of point clouds text pairs to train point clouds and text alignment. In this work, we propose a unified multimodal framework for open world 3D scene understanding, which is lightweight and scalable, while not requiring the production of additional point clouds text pairs.
3 Method
Figure 2: Overview of our proposed multi-modal unified learning framework, Uni-PLV. We utilize 2D foundation models to construct the regional semantic labels for images, which serve as the new classes supervisory signals. We input all category names as text prompts into the text encoder to get text embeddings. Given the spatial temporal aligned point clouds and multi-view images, we extract features using both 2D and 3D branches and compute similarities with text embeddings to obtain semantic categories for pixels and points. We distill classification logits and features of images onto the point cloud to shorten the feature distance of three modalities. Images and point clouds features are also input to vison-point matching module to predict whether points and pixels match. Simply providing text descriptions and point clouds, this framework could obtain 3D semantic understanding during inference.
3.1 Overview
@@
In contrast to previous works that achieved open vocabulary understanding by constructing 3D points-text pairs, our work transfers the open vocabulary capability from 2D to 3D by utilizing image regional semantic labels from 2D foundation models without any additional data processing (Sec.3.2). With the 2D-3D spatial mapping and pre-aligned images and texts, we design a multi-modal unified training framework that uses images as a bridge to embed point cloud features into the shared feature space of images and text. In Sec.3.3, we introduce main components of the proposed framework, the transformation of data flow, two knowledge distillation modules, and a vison-point matching module. Four task-specific optimization objectives are discussed in Sec.3.4. We introduce a multi-modal and multi-task training strategy to ensure stable and efficient training of point cloud and image branches (Sec.3.5). During inference, this framework only requires point clouds and class descriptions as input to compute feature similarity, selecting the most similar class as the semantic prediction for each point (Sec.3.6).
3.2 Image Data Processing
@@
Analysis. Due to the more accessible and larger image-text pairing datasets, recent 2D open-world methods have made significant progress by aligning images and text through contrastive learning. Despite the scarcity of point clouds and the complexity of constructing point clouds-text pairs, point clouds and images are usually captured together and spatially aligned through calibration, which inspires us to use images as a bridge to align point clouds and text into a unified feature space. Considering that image-level alignment data generated by large-scale visual language models such as CLIP could not satisfy the needs of point-level scene understanding, we utilize 2D foundation models to obtain pixel-level correspondences between text semantics and image regions. We would like to emphasize that, despite the predicted image region semantics being noisy and presenting numerous misalignments when transferred to the 3D scene through projection, as illustrated in Figure 1, our designed multi-modal unified training framework, employing images as a bridge, can accurately identify and locate point clouds for given categories.
@@
Region-Text Generation. Intuitively, we extract instance and pixel semantics from 2D vision-language foundation models. Specifically, given a set of images $\mathbf{I}=\{\mathbf{I}_{n}|n=1,...,N\}$ and the category list $\{t_{c}\}_{c=1}^{C}$ as input, where $t_{c}$ represents for the k-th category name, e.g., T = [”car”, ”pedestrain”, ”bicycle”, … , ”motorcycle”], the data pipeline output the bounding boxes $\mathbf{\hat{b}}$, instance mask $\mathbf{\hat{m}}$ and the category name $\mathbf{\hat{t}}$ for each image. We use GLEE [29] for instance mask and bounding box generation, which has been trained on large-scale datasets with high accuracy and generalization. We combine Grounding DINO [18] and SAM2 [25] to generate another set of instance labels. Bounding boxes were generated via the Grounding DINO, and each box was further segmented using SAM2 to produce instance masks. We obtain the region-pixel-text pairs $\langle\mathbf{\hat{b}},\mathbf{\hat{m}},\mathbf{\hat{t}}\rangle$, as well as the point clouds spatio-temporally aligned to the images, to train the designed 3D scene understanding network transferred from regional image-text alignment. 2D semantic labels for the experimental results in this paper are from GLEE, and related experiments on Grounding DINO with SAM2 can be found in the supporting materials.
3.3 Multi-Modal Alignment
3.3.1 Model Framework
@@
The proposed UniPLV consists of a frozen text encoder, image encoder-decoder, and point cloud segmentation branch, as illustrated in Figure 2. We input all category names as text prompts into the text encoder. We follow GLEE to apply global average pooling along the sequence dimension to get text embeddings $emb_{t}$. To support open-world understanding, we replace the classifier of image decoder and 3D segment head by the similarity measurement between perception features with text embeddings. Classification logits of 2D $L_{I}$ and 3D $L_{P}$ are computed as follows:
@@
At each iteration $e$ of the training, we have a set of images $\{\mathbf{I}_{e}^{k}\in\mathbb{R}^{H\times W\times 3}|k=1,...,K\}$ captured by $K$ cameras, where $H$ and $W$ are image dimensions. We also have the corresponding point cloud $\mathbf{P}_{e}\in\mathbb{R}^{N_{e}\times 3}$ with the mapping matrix $\Gamma_{lc}$ from LiDAR coordinate to image coordinate. Given the category list $\{t_{c}\}_{c=1}^{C}$, images $\{\mathbf{I}_{e}^{k}\}$ and the point cloud $\mathbf{P}_{e}$ as input, UniPLV could fine-tune the segmentation and detection of the image with the constructed regional image-text pairs $\langle\mathbf{\hat{b}},\mathbf{\hat{m}},\mathbf{\hat{t}}\rangle$, and provide the segmentation results of the point cloud corresponding to the given category list. The ultimate optimization goal of this framework is to embed point cloud features and image-text features into a unified feature space through multi-modal joint training, enabling the alignment of point cloud and text for open-world 3D scene understanding.
@@
3.3.2 Vision-Point Knowledge Distillation
To construct the framework in which image acts as a bridge for the joint embedding of point cloud features and pre-aligned image-text pairs into a unified feature space, we introduce two distillation modules: logit distillation and feature distillation.
Logit Distillation. The semantic classification logits $L_{I}$ of image pixels are obtained through similarity measurement between the image feature and all given categories txt embeddings as eq.1. Similarly, the semantic classification logits $L_{P}$ of point clouds are also obtained by calculating similarity with text as eq.2. We design the logit distillation $\mathcal{L}_{distill_{L}}$ supervising the point cloud classification logits $L_{P}$ with novel class semantics $S_{C_{n}}$ predicted by the image branch as follows:
$S_{C_{n}}=\sigma(L_{I})\to Proj(\Gamma_{lc},\mathbf{P}),\$ (3)

### v2

@@ Abstract
Open-world 3D scene understanding is a critical challenge that involves recognizing and distinguishing diverse objects and categories from 3D data, such as point clouds, without relying on manual annotations. Traditional methods struggle with this open-world task, especially due to the limitations of constructing extensive point cloud-text pairs and handling multimodal data effectively. In response to these challenges, we present UniPLV, a robust framework that unifies point clouds, images, and text within a single learning paradigm for comprehensive 3D scene understanding. UniPLV leverages images as a bridge to co-embed 3D points with pre-aligned images and text in a shared feature space, eliminating the need for labor-intensive point cloud-text pair crafting. Our framework achieves precise multimodal alignment through two innovative strategies: (i) logit and feature distillation modules between images and point clouds to enhance feature coherence, and (ii) a vision-point matching module that implicitly corrects 3D semantic predictions affected by projection inaccuracies from points to pixels. To further boost performance, we implement four task-specific losses alongside a two-stage training strategy. Extensive experiments demonstrate that UniPLV significantly surpasses state-of-the-art methods, with average improvements of 15.6% and 14.8% in semantic segmentation for Base-Annotated and Annotation-Free tasks, respectively. These results underscore UniPLV’s efficacy in pushing the boundaries of open-world 3D scene understanding. We will release the code to support future research and development.
I Introduction
Open-world 3D scene understanding is a fundamental yet challenging task that aims to enable models to recognize and distinguish open-set objects and categories directly from 3D data, such as point clouds, without relying on manual annotations. Unlike traditional closed-set approaches, which depend heavily on predefined categories and human-labeled datasets, open-world methods must generalize effectively to unseen objects and categories, a capability essential for real-world applications such as autonomous driving, robotics, and augmented or virtual reality. However, the inherent lack of annotations in large-scale point cloud data and the difficulty of integrating textual semantic knowledge make open-world 3D understanding a particularly demanding goal.
Figure 1: 3D semantics directly projected from 2D semantics suffer from significant noise. Without any 3D annotations, our method achieves strong performance in 3D open-world semantic understanding by jointly training and co-embedding 3D points with pre-aligned images and text.
Recent breakthroughs in Visual Language Models (VLM) have ushered in new possibilities for bridging vision and language modalities, paving the way for open-world semantic understanding across domains. Extensions of VLM to 3D tasks, such as frozen-VLM labeling approaches [1, 2], utilize pre-trained vision-language alignment models to transfer open-vocabulary semantics to 3D data, enabling open-world scene understanding without manually labeled 3D annotations. However, these methods tend to neglect region-level feature distillation and often rely on resource-heavy processes, including extensive feature extraction, hierarchical fusion, and large-volume storage [3]. Such inefficiencies limit their scalability to more advanced 3D networks and large-scale 3D scenes. On the other hand, pair-construction-based methods [4, 5, 6] attempt to align point clouds and text representations, much like VLM training, by generating point cloud-text pairs for contrastive learning. While methods like RegionPLC [6] achieve state-of-the-art performance, creating point cloud-text pairs is a labor-intensive process that lacks the scalability of large-scale text-image datasets easily obtainable from the Internet. This fundamental bottleneck leaves pair-construction-based methods unable to fully capitalize on the open-world generalization capabilities of VLMs.
Nonetheless, point clouds frequently appear alongside image modalities with spatial and temporal alignment, suggesting an alternative strategy: use 2D models for image semantic labeling, followed by projection into 3D space. This approach introduces potential errors due to calibration and motion artifacts, resulting in noise and omissions in semantic data, as shown in Fig. 1 (a). These challenges demand efficient methods to align text embeddings with point cloud features, minimizing projection inaccuracies. Given the spatial pairing of points and images, and the pre-existing image-text feature alignments in VLMs, images can act as bridges in a unified framework, easing the alignment of text with 3D point clouds.
Several critical challenges must be tackled to realize an open-world framework for 3D scene understanding with limited labeled 3D data: (i) co-embedding point cloud features with pre-aligned images and text into a shared feature space despite data constraints; (ii) addressing alignment errors caused by inaccurate projections between point clouds and images; and (iii) ensuring stable and effective multimodal training for robust performance across diverse scenes and categories.
To address these challenges, we propose UniPLV, a unified multimodal framework for open-world 3D scene understanding that eliminates the need for point cloud-text pair construction. UniPLV leverages pre-aligned visual language models and a point cloud understanding network, using images as a bridge to align point clouds and text in a shared feature space. To achieve this, we generate regional semantic labels for images using 2D foundation models and introduce logit and feature distillation techniques to align point cloud and image features. Additionally, we design a vision-point matching (VPM) module to refine point-pixel alignment and address projection errors. To tackle challenges such as gradient biases and long-tail distributions in multimodal training, we employ a two-stage training strategy with four task-specific losses: image-text alignment, point-text alignment, pixel-point matching, and distillation losses. Training begins with the image branch and transitions to joint training with balanced loss weights between modalities. During inference, UniPLV requires only text descriptions and raw point clouds to achieve robust 3D semantic understanding. Extensive experiments demonstrate UniPLV’s state-of-the-art performance on open-world semantic segmentation tasks using nuScenes and ScanNet, with further evaluations conducted on Waymo and SemanticKITTI.
In summary, our contributions are as follows:
@@ I Introduction
  • We propose a unified multimodal contrastive learning framework for open-world 3D scene understanding without requiring point cloud-text pair construction.
@@
  • We develop the vision-point matching module, logit distillation, and feature distillation to effectively co-embed point clouds, images, and text into a shared feature space.
@@
  • We introduce a two-stage optimization strategy and four task-specific weighted losses to ensure stable and efficient multimodal training.
  • Our framework achieves state-of-the-art performance on multiple open-world benchmarks, including nuScenes, ScanNet, Waymo, and SemanticKITTI.
II Related Works
@@
Open-Vocabulary 2D Scene Understanding. The recent advances of large visual language models have expanded the ability to understand 2D open-world scenes. There are two main directions: CLIP-based and Grounding. CLIP-based methods usually replace the linear projection feature with the CLIP feature and use CLIP for feature alignment, such as GLEE[7], DetCLIP series[8, 9, 10], RegionCLIP[11], OWL-ViT[12]. The input of the Grounding tasks is a picture and a corresponding description. Grounding tasks output object boxes at the corresponding positions in the image based on different descriptions [13, 14, 15, 16].
@@
Open-Vocabulary 3D Scene Understanding. Open-world 3D scene understanding aims to recognize objects that are not annotated. With the success of visual-language (VL) models, such as CLIP [17], numerous efforts [18, 19, 3] have emerged to transfer the VL knowledge to 3D scene understanding. Clip2Scene [18] utilizes frozen CLIP to obtain semantic labels of images, which are subsequently projected to guide a point cloud semantic segmentation. Further, Chen et al. [20] employ CLIP to pseudo-label 2D image pixels and transfer those labels to 3D points, using SAM to reduce noise. Guo et al. [21] use a frozen 2D network to map to 3D points obtained through Gaussian Splatting. OpenMask3D [19] employs a 3D instance segmentation network to create 3D masks, which are projected to obtain 2D masks. These 2D masks are then fed into CLIP to extract visual features and match them with textual features, ultimately obtaining 3D semantics. Since CLIP is trained on full-image and text alignments, its capability to comprehend specific regions is limited. OpenScene [2] facilitates point cloud and text alignment by projecting the prediction results from a frozen 2D vision model and employing distillation between image and point cloud features. However, OpenScene requires resource-intensive feature extraction and fusion, and the image backbone is fixed during training, making it unable to be easily expanded to more advanced 3D networks and 3D scenes. RegionPLC [22] and PLA [5] achieve open world 3D scene understanding by constructing a large number of point clouds text pairs to train point clouds and text alignment.
Figure 2: We introduce UniPLV, a multimodal learning framework that uses 2D foundation models to generate semantic labels for images as new class supervision. Category names are encoded as text prompts to create embeddings. For aligned point clouds and multi-view images, 2D and 3D branches extract features that are matched with text embeddings to categorize pixels and points. We distill image features to the point cloud to unify modalities and use a matching module to predict point-pixel correspondence. With just text and point clouds, Uni-PLV achieves effective 3D semantic understanding.
III Methodology
III-A Overview
@@
In contrast to previous works that achieved open vocabulary understanding by constructing 3D points-text pairs, our work transfers the open vocabulary capability from 2D to 3D by utilizing image regional semantic labels from 2D foundation models without any additional data processing. With the 2D-3D spatial mapping and pre-aligned images and texts, we design a multimodal unified training framework that uses images as a bridge to embed point cloud features into the shared feature space of images and text. We introduce the main components of the proposed framework, the transformation of data flow, two knowledge distillation modules, and a vision-point matching module. Four task-specific optimization objectives are integrated in training. We introduce a multimodal and multi-task training strategy to ensure stable and efficient training of point cloud and image branches. During inference, this framework only requires point clouds and class descriptions as input to compute feature similarity, selecting the most similar class as the semantic prediction for each point.
III-B Image Data Processing
@@
Analysis. Due to the more accessible and larger image-text pairing datasets, recent 2D open-world methods have made significant progress by aligning images and text through contrastive learning. Despite the scarcity of point clouds and the complexity of constructing point clouds-text pairs, point clouds and images are usually captured together and spatially aligned through calibration, which inspires us to use images as a bridge to align point clouds and text into a unified feature space. Considering that image-level alignment data generated by large-scale visual language models such as CLIP could not satisfy the needs of point-level scene understanding, we utilize 2D foundation models to obtain pixel-level correspondences between text semantics and image regions. We would like to emphasize that, despite the predicted image region semantics being noisy and presenting numerous misalignments when transferred to the 3D scene through projection, as illustrated in Fig. 1, our designed multimodal unified training framework, employing images as a bridge, can accurately identify and locate point clouds for given categories.
@@
Region-Text Generation. Intuitively, we extract instance and pixel semantics from 2D vision-language foundation models. Specifically, given a set of images $\mathbf{I}=\{\mathbf{I}_{n}|n=1,...,N\}$ and the category list $\{t_{c}\}_{c=1}^{C}$ as input, where $t_{c}$ represents for the k-th category name, e.g., T = [“car”, “pedestrain”, “bicycle”, …, “motorcycle”], the data pipeline output the bounding boxes $\mathbf{\hat{b}}$, instance mask $\mathbf{\hat{m}}$ and the category name $\mathbf{\hat{t}}$ for each image. We use GLEE [7] for instance mask and bounding box generation, which has been trained on large-scale datasets with high accuracy and generalization. We combine Grounding DINO [15] and SAM2 [23] to generate another set of instance labels. Bounding boxes were generated via the Grounding DINO, and each box was further segmented using SAM2 to produce instance masks. We obtain the region-pixel-text pairs $\langle\mathbf{\hat{b}},\mathbf{\hat{m}},\mathbf{\hat{t}}\rangle$, as well as the point clouds spatio-temporally aligned to the images, to train the designed 3D scene understanding network transferred from regional image-text alignment. 2D semantic labels for the experimental results in this paper are from GLEE, and related experiments on Grounding DINO with SAM2 can be found in the supporting materials.
III-C MultiModal Alignment
III-C1 Model Framework
@@
The proposed UniPLV consists of a frozen text encoder, image encoder-decoder, and point cloud segmentation branch, as illustrated in Fig. 2. We input all category names as text prompts into the text encoder. We follow GLEE to apply global average pooling along the sequence dimension to get text embeddings $emb_{t}$. To support open-world understanding, we replace the classifier of the image decoder and the 3D segment head with the similarity measurement between perception features with text embeddings. Classification logits of 2D $L_{I}$ and 3D $L_{P}$ are computed as follows:
@@
At each iteration $e$ of the training, we have a set of images $\{\mathbf{I}_{e}^{k}\in\mathbb{R}^{H\times W\times 3}|k=1,...,K\}$ captured by $K$ cameras, where $H$ and $W$ are image dimensions. We also have the corresponding point cloud $\mathbf{P}_{e}\in\mathbb{R}^{N_{e}\times 3}$ with the mapping matrix $\Gamma_{lc}$ from LiDAR coordinate to image coordinate. Given the category list $\{t_{c}\}_{c=1}^{C}$, images $\{\mathbf{I}_{e}^{k}\}$ and the point cloud $\mathbf{P}_{e}$ as input, UniPLV could fine-tune the segmentation and detection of the image with the constructed regional image-text pairs $\langle\mathbf{\hat{b}},\mathbf{\hat{m}},\mathbf{\hat{t}}\rangle$, and provide the segmentation results of the point cloud corresponding to the given category list. The ultimate optimization goal of this framework is to embed point cloud features and image-text features into a unified feature space through multimodal joint training, enabling the alignment of point cloud and text for open-world 3D scene understanding.
@@
III-C2 Vision-Point Knowledge Distillation
To construct the framework in which the image acts as a bridge for the joint embedding of point cloud features and pre-aligned image-text pairs into a unified feature space, we introduce two distillation modules: logit distillation and feature distillation.
Logit Distillation. The semantic classification logits $L_{I}$ of image pixels are obtained through similarity measurement between the image feature and all given categories’ txt embeddings as eq.1. Similarly, the semantic classification logits $L_{P}$ of point clouds are also obtained by calculating similarity with text as eq.2. We design the logit distillation $\mathcal{L}_{distill_{L}}$ supervising the point cloud classification logits $L_{P}$ with novel class semantics $S_{C_{n}}$ predicted by the image branch as follows:
$S_{C_{n}}=\sigma(L_{I})\to Proj(\Gamma_{lc},\mathbf{P}),\\$ (3)